Men arrive at a health centre and at random, at a constant rate of 0.2 per minute. Women
arrive at the same health centre independently and at random, at a constant rate of 0.3 per
minute.
(i) Find the probability that at least three women arrive at the health centre during a 5-
minute period. [4]
(ii) Find the probability that at least 2 men and at least 3 women arrive at the health
centre during a 5-minute period. [5]
1
Expert's answer
2020-04-01T10:42:15-0400

i) Let random variale "\xi" is the number of women arrived at the health centre during a 5-minute period. Then "\xi\in" Poisson ("\lambda") where "\lambda=5(0.3)=1.5".

We will find "P\{\xi=0\}, P\{\xi=1\}, P\{\xi=2\}" and calculate sum of these probabilities.

Using cumulative table we get "0.8088." Then our probability is "1-0.8088=0.1912."

ii) Let random variable "\eta" is the number of men arrived at the health centre during a 5-minute period. Then "\eta\in" Poisson ("\lambda") where "\lambda=5(0.2)=1."

We will find "P\{\xi=0\}, P\{\xi=1\}" and calculate sum of these probabilities.

Using cumulative table we get "0.7358". Then probability that at least 2 men arrive at the health centre during a 5-minute period is "1-0.7358=0.2642."

Then our probability is "(0.2642)(0.1912)\approx 0.05" (events are independent).